The perfectly competitive firm and market in the short run Consider a perfectly competitive market where demand is QD = 2,000 - 40P and quantity is measured in units while price is measured in dollars per unit. The long run supply is QS = 100P - 800.

a) Find the equilibrium price and the equilibrium quantity.

b) When the market is in equilibrium, what is the total expenditure in this market?

c) When the market is in equilibrium, what is the consumer surplus? What is the producer surplus?

d) Suppose the government introduces a quota that restricts quantity at Q = 800. How much in total consumer and producer surplus would be lost?

e) In a graph, measuring quantity in units along the horizontal axis and price in dollars per unit along the vertical axis illustrate your finding from part d).

f) Does the loss in consumer and producer surplus depends on whether the price settles at P =30, at P’ = 16, or any price in between? Discuss.

g) Call P the market price after the introduction of the quota and compute consumer surplus and producer surplus as a function of P. Using algebra, show that while CS and PS depend on P their sum does not.

Please answer all of them. Otherwise abstain from replying. Thanks
Answer #1

OD-2000-40P Os = 100 8-800 a) Em is achieved cords 200o-yop = pop-800 2800 = 1909 19-20 & 0=1200 3) pgtal Experten a 20x20o =